OpenStudy (anonymous):

what value of c makes the expression a perfect square? x^2+x+c

OpenStudy (anonymous):

1

myininaya (myininaya):

c=(1/2)^2=1/4 x^2+x+1/4=(x+1/2)^2

myininaya (myininaya):

x^2+bx+c to find the c that completes the square just take the number in front of x and divide by 2 and then square

myininaya (myininaya):

(b/2)^2

OpenStudy (anonymous):

@ myininaya what exactly is the answer?

myininaya (myininaya):

i said c=(1/2)^2=1/4

OpenStudy (anonymous):

ohhhh that makes soo much sense!

OpenStudy (anonymous):

thanks

myininaya (myininaya):

np